OpenStudy (anonymous):

1. Divide. (18x^3+12x^2-3x)/6x^2 2. Divide. (2x^3-x^2-24x+12)/(2x-1) 3. Simplify 2a+2/a^2-1 4. What is the undefined value of 3b^2+13b+4/b+4 5. Multiply. (x^2-x-42/x-8)(x^2-64/9x+54) 6.Divide (x^2+9+20/x^2-25) / (x+4/x-4) 7.Add. 8+ (x-1/x+4) 8. Subtract (x^2-1/x^2-x-2) - (x-1/x-2) 9. Solve (10/3x)+(4/3)=(7+x/2x) 10. Solve (x+12/x+4)=(x/x+8) 11. A group of college students is volunteering for Homes for the Community during their spring break. They are putting the finishing touches on a house they built. Working alone, Wade can paint a certain room in 7 hour

OpenStudy (anonymous):

5 dollars each

OpenStudy (anonymous):

I have you give you $5 per answer

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

11 seems incomplete?

OpenStudy (anonymous):

forget it man..

OpenStudy (anonymous):

okay 90 percent discount for u

OpenStudy (anonymous):

um no.

OpenStudy (anonymous):

100 percent discount?

OpenStudy (anonymous):

yeah you're not getting any of my money, you don't have to answer the questions if you dont want to

OpenStudy (anonymous):

i am just joking

OpenStudy (anonymous):

oh, sorry, talking to stragers online kinda freaks me out in the first place

OpenStudy (anonymous):

i got problem 11 dont worry about that one

OpenStudy (anonymous):

@passive13

OpenStudy (anonymous):

(4/b)) NUMBER 4

OpenStudy (anonymous):

2/a-1 NUMBER 3

OpenStudy (jack1):

number 7... you know how to add fractions...? eg 3/5 + 7/9 = ? so you find a common denominator (the bottom number), usually by multiplying the denominator of one by top and bottom of the other so: \[\frac{ 3}{ 5 } + \frac{ 7 }{ 9 } \rightarrow \frac{ 3\times 9 }{ 5\times9 } +\frac{ 7\times 5 }{ 9\times 5 } \] then you can add the top line if the denominators match \[\frac{ 27 }{ 45 } + \frac{ 35 }{ 45 } = \frac{ 27+35 }{ 45 } = \frac{ 62 }{ 45 }\] so use the same principal for Q7

OpenStudy (jack1):

so 8+ (x-1/x+4) is the same as: \[\frac{ 8 }{ 1 } + \frac{ (x-1) }{ (x+4) } = \frac{ 8\times(x+4) }{ 1\times(x+4) } + \frac{ 1\times(x-1) }{ 1(x+4) }\]